ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

exponentials 1991 MS Classification: not given Belinsky, Eduard, University of Zimbabwe, Zimbabwe Decomposition theorems and approximation by the “floating” system of exponentials The problem of approximation of function by polynomials with a prescribed \underline{number} of harmonics is considered. The iterational method can be applied in different functional spaces, as continuous functions, functions with uniformly convergent Fourier series, continuous analytic functions.